OpenStudy (jazzyluvsu123):

Medal and Fan !! 5-x over 2 + 4 = -6 Help plz!!

OpenStudy (jazzyluvsu123):

@agent0smith

OpenStudy (jazzyluvsu123):

@pooja195

OpenStudy (kgrendel0324):

Need help?

OpenStudy (jazzyluvsu123):

yes plz

OpenStudy (kgrendel0324):

Okay. I'm assuming you're solving for x?

OpenStudy (jazzyluvsu123):

yess

OpenStudy (agent0smith):

\[\large \frac{ 5-x }{ 2 }+ 4 = -6\]First subtract 4 from both sides

OpenStudy (kgrendel0324):

\[\frac{ 5-x }{ 2+4 }=-6\]

OpenStudy (kgrendel0324):

oh i see how you wrote it

OpenStudy (kgrendel0324):

5-x/2=-10

OpenStudy (jazzyluvsu123):